# The effect of accretion on scalar superradiant instability

### v1

@@ Appendix A Some useful formulae
$\alpha_{0^{\prime}}^{(011)}$ $=\left(\frac{2^{4}3^{2}}{\tau_{\mathrm{acc}}\mu}\right)^{{1}/{8}}\left[W_{0}\left(2^{{4}/{9}}3^{{2}/{3}}M_{\text{s,0}}^{-{4}/{3}}\tau_{\text{acc}}^{-{1}/{3}}\mu^{-{5}/{3}}e^{\frac{\alpha_{0}^{8}\tau_{\mathrm{acc}}\mu}{2^{4}3^{2}}}\right)\right]^{{1}/{8}},$ (90a)
$\alpha_{0^{\prime}}^{(022)}$ $=\left(\frac{3^{11}5^{2}}{2^{7}\tau_{\mathrm{acc}}\mu}\right)^{1/12}\left[W_{0}\left(2^{-17/5}3^{17/5}5^{-4/5}M_{\mathrm{s,0}}^{-6/5}\tau_{\mathrm{acc}}^{-1/5}\mu^{-7/5}e^{\frac{2^{7}\alpha_{0}^{12}\tau_{\mathrm{acc}}\mu}{3^{11}5^{2}}}\right)\right]^{1/12},$ (90b)
$\alpha_{0^{\prime}}^{(033)}$ $=\left(\frac{2^{17}5^{3}7^{2}}{3^{3}\tau_{\mathrm{acc}}\mu}\right)^{1/16}\left[W_{0}\left(2^{5/3}3^{-3/7}5^{3/7}7^{-6/7}M_{\mathrm{s,0}}^{-8/7}\tau_{\mathrm{acc}}^{-1/7}\mu^{-9/7}e^{\frac{3^{3}\alpha_{0}^{16}\tau_{\mathrm{acc}}\mu}{2^{17}5^{3}7^{2}}}\right)\right]^{1/16},$ (90c)
$\alpha_{0^{\prime}}^{(044)}$ $=\left(\frac{3^{8}5^{15}7^{3}}{2^{20}\tau_{\mathrm{acc}}\mu}\right)^{1/20}\left[W_{0}\left(2^{-140/27}3^{-2/9}5^{5/3}7^{1/3}M_{\mathrm{s,0}}^{-10/9}\tau_{\mathrm{acc}}^{-1/9}\mu^{-11/9}e^{\frac{2^{20}\alpha_{0}^{20}\tau_{\mathrm{acc}}\mu}{3^{8}5^{15}7^{3}}}\right)\right]^{1/20},$ (90d)
$\alpha_{0^{\prime}}^{(055)}$ $=\left(\frac{2^{11}3^{24}7^{3}11^{2}}{5^{7}\tau_{\mathrm{acc}}\mu}\right)^{1/24}\left[W_{0}\left(2^{3/11}3^{36/11}5^{-19/11}7^{3/11}11^{-10/11}M_{\mathrm{s,0}}^{-12/11}\tau_{\mathrm{acc}}^{-1/11}\mu^{-13/11}e^{\frac{5^{7}\alpha_{0}^{24}\tau_{\mathrm{acc}}\mu}{2^{11}3^{24}7^{3}11^{2}}}\right)\right]^{1/24},$ (90e)
$\alpha_{0^{\prime}}^{(066)}$ $=\left(\frac{5^{4}7^{19}11^{3}13^{2}}{2^{16}3^{2}\tau_{\mathrm{acc}}\mu}\right)^{1/28}\left[W_{0}\left(\frac{5^{4/13}7^{19/13}11^{3/13}}{2^{118/39}3^{2/13}13^{12/13}}M_{\mathrm{s,0}}^{-14/13}\tau_{\mathrm{acc}}^{-1/13}\mu^{-15/13}e^{\frac{2^{16}3^{2}\alpha_{0}^{28}\tau_{\mathrm{acc}}\mu}{5^{4}7^{19}11^{3}13^{2}}}\right)\right]^{1/28},$ (90f)
$\alpha_{0^{\prime}}^{(077)}$ $=\left(\frac{2^{50}3^{13}5^{6}11^{3}13^{3}}{7^{11}\tau_{\mathrm{acc}}\mu}\right)^{1/32}\left[W_{0}\left(\frac{2^{118/45}3^{13/15}11^{1/5}13^{1/5}}{5^{2/3}7^{9/5}}M_{\mathrm{s,0}}^{-16/15}\tau_{\mathrm{acc}}^{-1/15}\mu^{-17/15}e^{\frac{7^{11}\alpha_{0}^{32}\tau_{\mathrm{acc}}\mu}{2^{50}3^{13}5^{6}11^{3}13^{3}}}\right)\right]^{1/32}.$ (90g)
@@
$\alpha_{1}^{(011)}$ $=\left[-\frac{2^{4}3}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{4}\tau_{\mathrm{acc}}\mu^{5}a_{*0^{\prime}}^{-4}}{2^{4}3}e^{-\frac{\alpha_{0^{\prime}}^{8}\tau_{\mathrm{acc}}\mu}{2^{4}3}}\right)\right]^{1/8},$ (91)
$\alpha_{1}^{(022)}$ $=\left[-\frac{2^{-7}3^{11}5}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{6}\tau_{\mathrm{acc}}\mu^{7}a_{*0^{\prime}}^{-6}}{2^{-13}3^{11}5}e^{-\frac{\alpha_{0^{\prime}}^{12}\tau_{\mathrm{acc}}\mu}{2^{-7}3^{11}5}}\right)\right]^{1/12},$ (92)
$\alpha_{1}^{(033)}$ $=\left[-\frac{2^{17}3^{-3}5^{3}7}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{8}\tau_{\mathrm{acc}}\mu^{9}a_{*0^{\prime}}^{-8}}{2^{17}3^{-11}5^{3}7}e^{-\frac{\alpha_{0^{\prime}}^{16}\tau_{\mathrm{acc}}\mu}{2^{17}3^{-3}5^{3}7}}\right)\right]^{1/16},$ (93)
$\alpha_{1}^{(044)}$ $=\left[-\frac{2^{-20}3^{6}5^{15}7^{3}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{10}\tau_{\mathrm{acc}}\mu^{11}a_{*0^{\prime}}^{-10}}{2^{-40}3^{6}5^{15}7^{3}}e^{-\frac{\alpha_{0^{\prime}}^{20}\tau_{\mathrm{acc}}\mu}{2^{-20}3^{6}5^{15}7^{3}}}\right)\right]^{1/20},$ (94)
$\alpha_{1}^{(055)}$ $=\left[-\frac{2^{11}3^{24}5^{-7}7^{3}11}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{12}\tau_{\mathrm{acc}}\mu^{13}a_{*0^{\prime}}^{-12}}{2^{11}3^{24}5^{-19}7^{3}11}e^{-\frac{\alpha_{0^{\prime}}^{24}\tau_{\mathrm{acc}}\mu}{2^{11}3^{24}5^{-7}7^{3}11}}\right)\right]^{1/24},$ (95)
$\alpha_{1}^{(066)}$ $=\left[-\frac{2^{-16}3^{-2}5^{4}7^{19}11^{3}13}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{14}\tau_{\mathrm{acc}}\mu^{15}a_{*0^{\prime}}^{-14}}{2^{-30}3^{-16}5^{4}7^{19}11^{3}13}e^{-\frac{\alpha_{0^{\prime}}^{28}\tau_{\mathrm{acc}}\mu}{2^{-16}3^{-2}5^{4}7^{19}11^{3}13}}\right)\right]^{1/28},$ (96)
$\alpha_{1}^{(077)}$ $=\left[-\frac{2^{50}3^{12}5^{5}7^{-11}11^{3}13^{3}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{16}\tau_{\mathrm{acc}}\mu^{17}a_{*0^{\prime}}^{-16}}{2^{50}3^{12}5^{5}7^{-27}11^{3}13^{3}}e^{-\frac{\alpha_{0^{\prime}}^{32}\tau_{\mathrm{acc}}\mu}{2^{50}3^{12}5^{5}7^{-11}11^{3}13^{3}}}\right)\right]^{1/32}.$ (97)
@@
$\dot{E}_{\mathrm{SR}}^{(0ll)}=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3}{l}\sqrt{\frac{3}{2}}-\frac{31}{2l^{2}}\alpha+\mathcal{O}(\alpha^{2})\right].$ (98)
@@
$\dot{E}_{\mathrm{SR}}^{(011)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[3\sqrt{\frac{3}{2}}-\frac{31\alpha}{2}+\frac{787\alpha^{2}}{8\sqrt{6}}-\frac{20305\alpha^{3}}{216}+\frac{1564549\alpha^{4}}{3456\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99a)
$\dot{E}_{\mathrm{SR}}^{(022)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{2}-\frac{31\alpha}{8}+\frac{65}{16}\sqrt{\frac{3}{2}}\alpha^{2}-\frac{9743\alpha^{3}}{1728}+\frac{223999\alpha^{4}}{17280\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99b)
$\dot{E}_{\mathrm{SR}}^{(033)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\sqrt{\frac{3}{2}}-\frac{31\alpha}{18}+\frac{3103\alpha^{2}}{864\sqrt{6}}-\frac{75955\alpha^{3}}{69984}+\frac{151571899\alpha^{4}}{94058496\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99c)
$\dot{E}_{\mathrm{SR}}^{(044)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{4}-\frac{31\alpha}{32}+\frac{4831\alpha^{2}}{3200\sqrt{6}}-\frac{233279\alpha^{3}}{691200}+\frac{5079739\alpha^{4}}{13824000\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99d)
$\dot{E}_{\mathrm{SR}}^{(055)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{5}-\frac{31\alpha}{50}+\frac{257\sqrt{\frac{3}{2}}\alpha^{2}}{1000}-\frac{18433\alpha^{3}}{135000}+\frac{13894871\alpha^{4}}{118800000\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99e)
$\dot{E}_{\mathrm{SR}}^{(066)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{\sqrt{\frac{3}{2}}}{2}-\frac{31\alpha}{72}+\frac{9427\alpha^{2}}{21168\sqrt{6}}-\frac{447455\alpha^{3}}{6858432}+\frac{1205216951\alpha^{4}}{26212927104\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (99f)
$\dot{E}_{\mathrm{SR}}^{(077)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{7}-\frac{31\alpha}{98}+\frac{12295\alpha^{2}}{43904\sqrt{6}}-\frac{290131\alpha^{3}}{8297856}+\frac{1554440081\alpha^{4}}{74348789760\sqrt{6}}+\mathcal{O}(\alpha^{5})\right].$ (99g)
@@
${M}_{\mathrm{s,Regge}}^{(0ll)}=\frac{\alpha^{2}}{\alpha_{0}}\left[\frac{3}{2l}\sqrt{\frac{3}{2}}-\frac{2}{3l^{2}}\alpha+\mathcal{O}(\alpha^{2})\right].$ (100)
@@
${M}_{\mathrm{s,Regge}}^{(011)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{3}-\frac{473\alpha^{2}}{32\sqrt{6}}-\frac{223\alpha^{3}}{270}+\frac{999421\alpha^{4}}{20736\sqrt{6}}\right),$ (101a)
${M}_{\mathrm{s,Regge}}^{(022)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{4}\sqrt{\frac{3}{2}}-\frac{\alpha}{6}-\frac{5}{8}\sqrt{\frac{3}{2}}\alpha^{2}-\frac{\alpha^{3}}{27}+\frac{1313\alpha^{4}}{810\sqrt{6}}\right),$ (101b)
${M}_{\mathrm{s,Regge}}^{(033)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{1}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{27}-\frac{1937\alpha^{2}}{3456\sqrt{6}}-\frac{487\alpha^{3}}{87480}+\frac{125563483\alpha^{4}}{564350976\sqrt{6}}\right),$ (101c)
${M}_{\mathrm{s,Regge}}^{(044)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{8}\sqrt{\frac{3}{2}}-\frac{\alpha}{24}-\frac{761\alpha^{2}}{3200\sqrt{6}}-\frac{151\alpha^{3}}{108000}+\frac{561893\alpha^{4}}{10368000\sqrt{6}}\right),$ (101d)
${M}_{\mathrm{s,Regge}}^{(055)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{10}\sqrt{\frac{3}{2}}-\frac{2\alpha}{75}-\frac{163\sqrt{\frac{3}{2}}\alpha^{2}}{4000}-\frac{79\alpha^{3}}{168750}+\frac{2577331\alpha^{4}}{142560000\sqrt{6}}\right),$ (101e)
${M}_{\mathrm{s,Regge}}^{(066)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{1}{4}\sqrt{\frac{3}{2}}-\frac{\alpha}{54}-\frac{751\alpha^{2}}{10584\sqrt{6}}-\frac{101\alpha^{3}}{535815}+\frac{9044053\alpha^{4}}{1228730958\sqrt{6}}\right),$ (101f)
${M}_{\mathrm{s,Regge}}^{(077)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{14}\sqrt{\frac{3}{2}}-\frac{2\alpha}{147}-\frac{7865\alpha^{2}}{175616\sqrt{6}}-\frac{179\alpha^{3}}{2074464}+\frac{1534182161\alpha^{4}}{446092738560\sqrt{6}}\right).$ (101g)

### v2

@@ Appendix A Some useful formulae
$\alpha_{0^{\prime}}^{(011)}$ $=\left(\frac{2^{4}3^{2}}{\tau_{\mathrm{acc}}\mu}\right)^{{1}/{8}}\left[W_{0}\left(2^{{4}/{9}}3^{{2}/{3}}M_{\text{s,0}}^{-{4}/{3}}\tau_{\text{acc}}^{-{1}/{3}}\mu^{-{5}/{3}}e^{\frac{\alpha_{0}^{8}\tau_{\mathrm{acc}}\mu}{2^{4}3^{2}}}\right)\right]^{{1}/{8}},$ (A.13a)
$\alpha_{0^{\prime}}^{(022)}$ $=\left(\frac{3^{11}5^{2}}{2^{7}\tau_{\mathrm{acc}}\mu}\right)^{1/12}\left[W_{0}\left(2^{-17/5}3^{17/5}5^{-4/5}M_{\mathrm{s,0}}^{-6/5}\tau_{\mathrm{acc}}^{-1/5}\mu^{-7/5}e^{\frac{2^{7}\alpha_{0}^{12}\tau_{\mathrm{acc}}\mu}{3^{11}5^{2}}}\right)\right]^{1/12},$ (A.13b)
$\alpha_{0^{\prime}}^{(033)}$ $=\left(\frac{2^{17}5^{3}7^{2}}{3^{3}\tau_{\mathrm{acc}}\mu}\right)^{1/16}\left[W_{0}\left(2^{5/3}3^{-3/7}5^{3/7}7^{-6/7}M_{\mathrm{s,0}}^{-8/7}\tau_{\mathrm{acc}}^{-1/7}\mu^{-9/7}e^{\frac{3^{3}\alpha_{0}^{16}\tau_{\mathrm{acc}}\mu}{2^{17}5^{3}7^{2}}}\right)\right]^{1/16},$ (A.13c)
$\begin{split}\alpha_{0^{\prime}}^{(044)}&=\left(\frac{3^{8}5^{15}7^{3}}{2^{20}\tau_{\mathrm{acc}}\mu}\right)^{1/20}\\ &\hskip 14.22636pt\times\left[W_{0}\left(2^{-140/27}3^{-2/9}5^{5/3}7^{1/3}M_{\mathrm{s,0}}^{-10/9}\tau_{\mathrm{acc}}^{-1/9}\mu^{-11/9}e^{\frac{2^{20}\alpha_{0}^{20}\tau_{\mathrm{acc}}\mu}{3^{8}5^{15}7^{3}}}\right)\right]^{1/20},\end{split}$ (A.13d)
$\begin{split}\alpha_{0^{\prime}}^{(055)}&=\left(\frac{2^{11}3^{24}7^{3}11^{2}}{5^{7}\tau_{\mathrm{acc}}\mu}\right)^{1/24}\\ &\hskip 14.22636pt\times\left[W_{0}\left(2^{3/11}3^{36/11}5^{-19/11}7^{3/11}11^{-10/11}M_{\mathrm{s,0}}^{-12/11}\tau_{\mathrm{acc}}^{-1/11}\mu^{-13/11}e^{\frac{5^{7}\alpha_{0}^{24}\tau_{\mathrm{acc}}\mu}{2^{11}3^{24}7^{3}11^{2}}}\right)\right]^{1/24},\end{split}$ (A.13e)
$\begin{split}\alpha_{0^{\prime}}^{(066)}&=\left(\frac{5^{4}7^{19}11^{3}13^{2}}{2^{16}3^{2}\tau_{\mathrm{acc}}\mu}\right)^{1/28}\\ &\hskip 14.22636pt\times\left[W_{0}\left(\frac{5^{4/13}7^{19/13}11^{3/13}}{2^{118/39}3^{2/13}13^{12/13}}M_{\mathrm{s,0}}^{-14/13}\tau_{\mathrm{acc}}^{-1/13}\mu^{-15/13}e^{\frac{2^{16}3^{2}\alpha_{0}^{28}\tau_{\mathrm{acc}}\mu}{5^{4}7^{19}11^{3}13^{2}}}\right)\right]^{1/28},\end{split}$ (A.13f)
$\begin{split}\alpha_{0^{\prime}}^{(077)}&=\left(\frac{2^{50}3^{13}5^{6}11^{3}13^{3}}{7^{11}\tau_{\mathrm{acc}}\mu}\right)^{1/32}\\ &\hskip 14.22636pt\times\left[W_{0}\left(\frac{2^{118/45}3^{13/15}11^{1/5}13^{1/5}}{5^{2/3}7^{9/5}}M_{\mathrm{s,0}}^{-16/15}\tau_{\mathrm{acc}}^{-1/15}\mu^{-17/15}e^{\frac{7^{11}\alpha_{0}^{32}\tau_{\mathrm{acc}}\mu}{2^{50}3^{13}5^{6}11^{3}13^{3}}}\right)\right]^{1/32}.\end{split}$ (A.13g)
@@
$\alpha_{1}^{(011)}$ $=\left[-\frac{2^{4}3}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{4}\tau_{\mathrm{acc}}\mu^{5}a_{*0^{\prime}}^{-4}}{2^{4}3}e^{-\frac{\alpha_{0^{\prime}}^{8}\tau_{\mathrm{acc}}\mu}{2^{4}3}}\right)\right]^{1/8},$ (A.14a)
$\alpha_{1}^{(022)}$ $=\left[-\frac{2^{-7}3^{11}5}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{6}\tau_{\mathrm{acc}}\mu^{7}a_{*0^{\prime}}^{-6}}{2^{-13}3^{11}5}e^{-\frac{\alpha_{0^{\prime}}^{12}\tau_{\mathrm{acc}}\mu}{2^{-7}3^{11}5}}\right)\right]^{1/12},$ (A.14b)
$\alpha_{1}^{(033)}$ $=\left[-\frac{2^{17}3^{-3}5^{3}7}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{8}\tau_{\mathrm{acc}}\mu^{9}a_{*0^{\prime}}^{-8}}{2^{17}3^{-11}5^{3}7}e^{-\frac{\alpha_{0^{\prime}}^{16}\tau_{\mathrm{acc}}\mu}{2^{17}3^{-3}5^{3}7}}\right)\right]^{1/16},$ (A.14c)
$\alpha_{1}^{(044)}$ $=\left[-\frac{2^{-20}3^{6}5^{15}7^{3}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{10}\tau_{\mathrm{acc}}\mu^{11}a_{*0^{\prime}}^{-10}}{2^{-40}3^{6}5^{15}7^{3}}e^{-\frac{\alpha_{0^{\prime}}^{20}\tau_{\mathrm{acc}}\mu}{2^{-20}3^{6}5^{15}7^{3}}}\right)\right]^{1/20},$ (A.14d)
$\alpha_{1}^{(055)}$ $=\left[-\frac{2^{11}3^{24}5^{-7}7^{3}11}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{12}\tau_{\mathrm{acc}}\mu^{13}a_{*0^{\prime}}^{-12}}{2^{11}3^{24}5^{-19}7^{3}11}e^{-\frac{\alpha_{0^{\prime}}^{24}\tau_{\mathrm{acc}}\mu}{2^{11}3^{24}5^{-7}7^{3}11}}\right)\right]^{1/24},$ (A.14e)
$\alpha_{1}^{(066)}$ $=\left[-\frac{2^{-16}3^{-2}5^{4}7^{19}11^{3}13}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{14}\tau_{\mathrm{acc}}\mu^{15}a_{*0^{\prime}}^{-14}}{2^{-30}3^{-16}5^{4}7^{19}11^{3}13}e^{-\frac{\alpha_{0^{\prime}}^{28}\tau_{\mathrm{acc}}\mu}{2^{-16}3^{-2}5^{4}7^{19}11^{3}13}}\right)\right]^{1/28},$ (A.14f)
$\alpha_{1}^{(077)}$ $=\left[-\frac{2^{50}3^{12}5^{5}7^{-11}11^{3}13^{3}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left(-\frac{M_{\mathrm{s0^{\prime}}}^{16}\tau_{\mathrm{acc}}\mu^{17}a_{*0^{\prime}}^{-16}}{2^{50}3^{12}5^{5}7^{-27}11^{3}13^{3}}e^{-\frac{\alpha_{0^{\prime}}^{32}\tau_{\mathrm{acc}}\mu}{2^{50}3^{12}5^{5}7^{-11}11^{3}13^{3}}}\right)\right]^{1/32}.$ (A.14g)
@@
$\dot{E}_{\mathrm{SR}}^{(0ll)}=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3}{l}\sqrt{\frac{3}{2}}-\frac{31}{2l^{2}}\alpha+\mathcal{O}(\alpha^{2})\right].$ (A.15)
@@
$\dot{E}_{\mathrm{SR}}^{(011)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[3\sqrt{\frac{3}{2}}-\frac{31\alpha}{2}+\frac{787\alpha^{2}}{8\sqrt{6}}-\frac{20305\alpha^{3}}{216}+\frac{1564549\alpha^{4}}{3456\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16a)
$\dot{E}_{\mathrm{SR}}^{(022)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{2}-\frac{31\alpha}{8}+\frac{65}{16}\sqrt{\frac{3}{2}}\alpha^{2}-\frac{9743\alpha^{3}}{1728}+\frac{223999\alpha^{4}}{17280\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16b)
$\dot{E}_{\mathrm{SR}}^{(033)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\sqrt{\frac{3}{2}}-\frac{31\alpha}{18}+\frac{3103\alpha^{2}}{864\sqrt{6}}-\frac{75955\alpha^{3}}{69984}+\frac{151571899\alpha^{4}}{94058496\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16c)
$\dot{E}_{\mathrm{SR}}^{(044)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{4}-\frac{31\alpha}{32}+\frac{4831\alpha^{2}}{3200\sqrt{6}}-\frac{233279\alpha^{3}}{691200}+\frac{5079739\alpha^{4}}{13824000\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16d)
$\dot{E}_{\mathrm{SR}}^{(055)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{5}-\frac{31\alpha}{50}+\frac{257\sqrt{\frac{3}{2}}\alpha^{2}}{1000}-\frac{18433\alpha^{3}}{135000}+\frac{13894871\alpha^{4}}{118800000\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16e)
$\dot{E}_{\mathrm{SR}}^{(066)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{\sqrt{\frac{3}{2}}}{2}-\frac{31\alpha}{72}+\frac{9427\alpha^{2}}{21168\sqrt{6}}-\frac{447455\alpha^{3}}{6858432}+\frac{1205216951\alpha^{4}}{26212927104\sqrt{6}}+\mathcal{O}(\alpha^{5})\right],$ (A.16f)
$\dot{E}_{\mathrm{SR}}^{(077)}$ $=\frac{\alpha M}{{\tau}_{\mathrm{acc}}}\left[\frac{3\sqrt{\frac{3}{2}}}{7}-\frac{31\alpha}{98}+\frac{12295\alpha^{2}}{43904\sqrt{6}}-\frac{290131\alpha^{3}}{8297856}+\frac{1554440081\alpha^{4}}{74348789760\sqrt{6}}+\mathcal{O}(\alpha^{5})\right].$ (A.16g)
@@
${M}_{\mathrm{s,Regge}}^{(0ll)}=\frac{\alpha^{2}}{\alpha_{0}}\left[\frac{3}{2l}\sqrt{\frac{3}{2}}-\frac{2}{3l^{2}}\alpha+\mathcal{O}(\alpha^{2})\right].$ (A.17)
@@
${M}_{\mathrm{s,Regge}}^{(011)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{3}-\frac{473\alpha^{2}}{32\sqrt{6}}-\frac{223\alpha^{3}}{270}+\frac{999421\alpha^{4}}{20736\sqrt{6}}\right),$ (A.18a)
${M}_{\mathrm{s,Regge}}^{(022)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{4}\sqrt{\frac{3}{2}}-\frac{\alpha}{6}-\frac{5}{8}\sqrt{\frac{3}{2}}\alpha^{2}-\frac{\alpha^{3}}{27}+\frac{1313\alpha^{4}}{810\sqrt{6}}\right),$ (A.18b)
${M}_{\mathrm{s,Regge}}^{(033)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{1}{2}\sqrt{\frac{3}{2}}-\frac{2\alpha}{27}-\frac{1937\alpha^{2}}{3456\sqrt{6}}-\frac{487\alpha^{3}}{87480}+\frac{125563483\alpha^{4}}{564350976\sqrt{6}}\right),$ (A.18c)
${M}_{\mathrm{s,Regge}}^{(044)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{8}\sqrt{\frac{3}{2}}-\frac{\alpha}{24}-\frac{761\alpha^{2}}{3200\sqrt{6}}-\frac{151\alpha^{3}}{108000}+\frac{561893\alpha^{4}}{10368000\sqrt{6}}\right),$ (A.18d)
${M}_{\mathrm{s,Regge}}^{(055)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{10}\sqrt{\frac{3}{2}}-\frac{2\alpha}{75}-\frac{163\sqrt{\frac{3}{2}}\alpha^{2}}{4000}-\frac{79\alpha^{3}}{168750}+\frac{2577331\alpha^{4}}{142560000\sqrt{6}}\right),$ (A.18e)
${M}_{\mathrm{s,Regge}}^{(066)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{1}{4}\sqrt{\frac{3}{2}}-\frac{\alpha}{54}-\frac{751\alpha^{2}}{10584\sqrt{6}}-\frac{101\alpha^{3}}{535815}+\frac{9044053\alpha^{4}}{1228730958\sqrt{6}}\right),$ (A.18f)
${M}_{\mathrm{s,Regge}}^{(077)}$ $=\frac{\alpha^{2}}{\alpha_{0}}\left(\frac{3}{14}\sqrt{\frac{3}{2}}-\frac{2\alpha}{147}-\frac{7865\alpha^{2}}{175616\sqrt{6}}-\frac{179\alpha^{3}}{2074464}+\frac{1534182161\alpha^{4}}{446092738560\sqrt{6}}\right).$ (A.18g)